The heigth of a person is 1.65 m. Express it into cm and mm 
 

Grade:6

4 Answers

Arun
4 years ago
Dear student
 
1m = 100 cm
And 1 m = 1000 mm
 
Hemce 1.65 m = 165 cm
 
And 1.65 m = 1650 mm
 
Hope it helps
 
Regards
Arun
Vikas TU
4 years ago
Dear student 
For converting meter in to cm multiply with 100 and for converting m to mm multiply it with 1000. 
Hope his clears now. 
Good Luck 
Yashraj
3 years ago
length in cm is 165cm and 1650mm 
…..........................................
1m=100cm
1cm=10mm
…............................................
KALYAN
3 years ago
Dear Aman Suthar,
for converting meters to centimeters and meters to millimeter we need to know unit coversions
 
1 meter = 100 centi meter
1 centimeter = 10 milli meter
1 meter = 100cm*10mm=1000 milli meter
so,
1.65m= 1.65*100 = 165cm
1.65m= 2.65*1000 = 1650 mm